Emmanuel Zacchini launched as a human cannonball in 1940 soared over three Ferris wheels each 18.0 m high covering a horizontal distance of 10.63m. a.) Assuming that the point of projection is 2.50 m above the ground and that he landed safely on a net placed at the same level, find his initial velocity. b.) Assuming that he missed the net and landed on the ground, with what velocity will he strike the ground?


1
Expert's answer
2021-11-07T19:25:17-0500

We have range and height which can give us the tangent of the angle of launch. Note: the height we're going to use here is the height between the highest point of trajectory and the 2.5 m where the launch was initiated:


"H=\frac{v^2\sin^2\theta}{2g}=\frac{v^2·\sin\theta·\sin\theta}{2g},\\\space\\
R=\frac{v^2\sin(2\theta)}{g}=\frac{v^2·2\sin\theta·\cos\theta}{g},\\\space\\

\frac HR=\frac14\tan\theta,\\\space\\
\theta=\arctan\frac{4H}{R}=\arctan\frac{4(18-2.5)}{10.63}=80°.\\\space\\
v=\sqrt\frac{Rg}{\sin(2\theta)}=17.5\text{ m/s}."

Assuming that he missed the net and landed on the ground:


"u_y=\sqrt{2g(H+h)}=18.8\text{ m/s},\\
u_x=v_x=v\cos\theta=3.04\text{ m/s},\\
u=\sqrt{u_x^2+u_y^2}=19\text{ m/s}."